OpenStudy (anonymous):

please help. How would i convert f(x) = log (x) + 4 using the change of base formula?

OpenStudy (phi):

Do they give any more background? convert f(x) to what?

OpenStudy (anonymous):

i needs to graph it, but in order to put this function on geogebra i need to convert it

OpenStudy (phi):

what base is it in now?

OpenStudy (anonymous):

it worked?

OpenStudy (phi):

yes

OpenStudy (anonymous):

oh, thanks!

OpenStudy (phi):

OK, grapher is a different program, but in geogebra I found lg(x) means log base 10 and ln(x) means log base e

OpenStudy (phi):

I started geogebra at the bottom I typed in y= lg(x)+4 and then return. It plotted the curve

OpenStudy (anonymous):

how will i know which to use? in or lg for diff problems?

OpenStudy (phi):

I think if you see log(x) that generally means log base 10 so use lg(x) in geogebra ln(x) means natural logarithm (log base e)

OpenStudy (phi):

it is little L little n in ln (log natural )

OpenStudy (anonymous):

how would i write this one? log0.2 (x + 6)

OpenStudy (anonymous):

i typed this in, y=lg0.2(x+6) and it didnt work

OpenStudy (phi):

Is that supposed to be log base 0.2 ?

OpenStudy (anonymous):

whoops, yes, does that mean i type in?

OpenStudy (anonymous):

y=in0.2(x+6) ?

OpenStudy (phi):

you have to use log_ugly_base(x) = lg(x)/lg( ugly_base)

OpenStudy (phi):

so you want y= lg(x+6)/lg(0.2)

OpenStudy (phi):

replace \[ \log_{0.2}(x) = \lg(x)/\lg(0.2) \]

OpenStudy (phi):

Did you get it to work?

OpenStudy (anonymous):

yes. you´re the beez kneez. thanks!

OpenStudy (phi):

you could also use y= ln(x+6)/ln(0.2) it should give the same plot